# Resolving Spurious Multifractality in Discrete Systems: A Finite-Size Scaling Protocol for MFDFA in the 2D Ising Model

### 04 Mar 2026 v1

@@ Abstract
Multifractal Detrended Fluctuation Analysis (MFDFA) has emerged as a standard tool for characterizing scale invariance in complex systems, yet its application to discrete spin models is frequently marred by reports of “spurious multifractality” that contradict established theory. In this work, we resolve this controversy by establishing a rigorous protocol for the analysis of discrete lattice snapshots. Using the 2D Ising model as a benchmark, we demonstrate that the previously reported broad singularity spectra [1] are finite-size artifacts dominated by lattice discreteness effects in the negative moment regime ($q<0$). By restricting the analysis to positive moments and performing a systematic Finite-Size Scaling (FSS) analysis, we show that the spectral width collapses to zero ($\Delta\alpha\to 0$) in the thermodynamic limit. The method accurately recovers the monofractal exponent of the Ising universality class ($\alpha\approx H\approx 0.875$), consistent with Conformal Field Theory. To validate the discriminatory power of this protocol, we contrast these findings with the Random Bond Ising Model (RBIM), showing that quenched disorder induces a genuine, broad multifractal spectrum ($\Delta\alpha\approx 0.23$) that survives scaling. Furthermore, we propose a theoretical interpretation where the MFDFA polynomial detrending functions as a phenomenological Renormalization Group filter, suppressing analytic background fields (irrelevant operators) to isolate the singular critical behavior. These results define a robust methodology for distinguishing between clean and disorder-dominated criticality in finite systems.
@@ I Introduction
A survey of the recent literature reveals a persistent anomaly: numerous numerical studies report broad multifractal spectra in the pure 2D Ising model. These findings imply the existence of a hierarchy of scaling exponents, a result that contradicts the unique scaling dimension mandated by the RG fixed point. This discrepancy—often termed ”spurious multifractality”—raises a fundamental question: Does the reported multifractality reflect hidden complexity in the Ising phase transition, or does it signal a breakdown of the MFDFA methodology when applied to discrete variables in finite systems?
@@
To validate that this collapse to monofractality is a correct physical result rather than a loss of methodological sensitivity, we contrast our findings with the Random Bond Ising Model (RBIM). The RBIM, characterized by quenched bond disorder, is known to exhibit ”Griffiths phases” [6] —rare regions that order locally at temperatures different from the bulk $T_{c}$. We demonstrate that our MFDFA protocol successfully detects a genuine, broad multifractal spectrum ($\Delta\alpha\approx 0.23$) in the disordered case that survives finite-size scaling. This establishes the method as a robust discriminator between ”clean” criticality and ”dirty” multifractality.
@@ III.1 Model Hamiltonians and Monte Carlo Scheme
The equilibrium configurations were generated using the Metropolis single-spin flip algorithm. To accommodate the critical slowing down near $T_{c}$, the number of Monte Carlo steps (MCS) per site was dynamically adjusted with system size, typically scaling as $L^{z}$ with $z\approx 2.17$. We simulated lattice sizes ranging from $L=32$ to $L=256$. For each temperature and system size, we collected an ensemble of $10^{4}$ independent equilibrium snapshots to ensure statistical convergence of the high-order moments.
@@ 4. Scaling and Singularity Spectrum:
The MFDFA implementation employed here follows the standard algorithm established by Kantelhardt et al. [4], with optimizations for computational efficiency when analyzing large ensembles of lattice configurations. Our research group has successfully applied MFDFA and related multifractal methods to diverse systems including climate indices [8], financial time series [9], and biological sequences [10], establishing the methodological foundation for the present investigation of discrete lattice models.
@@ IV.1 The Critical Fixed Point: Restoration of Monofractality
We first examine the scaling behavior of the pure 2D Ising model. A prerequisite for any valid multifractal analysis is the accurate recovery of the known Hurst exponent at the critical point. Figure 2 displays the temperature dependence of the second-order Hurst exponent $H(q=2)$ for a large lattice ($L=256$). As the system approaches the critical temperature $T_{c}\approx 2.269$, the exponent exhibits a sharp crossover, intersecting the theoretical prediction $H=0.875$ precisely at the transition point. This agreement confirms that the spatial MFDFA protocol correctly captures the anomalous dimension of the magnetization field, satisfying the scaling relation $H=1-\eta/2$.
@@
Figure 2: Evolution of the Hurst exponent $H$ ($q=2$) vs Temperature for $L=256$. The intersection with the theoretical value $H=0.875$ (blue dotted line) occurs precisely at the critical temperature $T_{c}$ (red dashed line).
@@
The central result of this study concerns the width of the singularity spectrum $f(\alpha)$. Previous studies reporting ”spurious multifractality” typically observed a broad spectral width $\Delta\alpha=\alpha_{max}-\alpha_{min}>0$. In Figure 3, we present the evolution of the singularity spectrum for system sizes ranging from $L=32$ to $L=256$, restricted to the positive moment regime ($q>0$). For small system sizes ($L=32$), the spectrum indeed appears broad, mimicking a multifractal structure. However, this width is not intrinsic. As $L$ increases, we observe a dramatic narrowing of the spectrum. The distinct branches of the $f(\alpha)$ curve progressively collapse onto a single vertical line centered at $\alpha\approx 0.875$.
Figure 3: Evolution of the singularity spectra $f(\alpha)$ for $q>0$. Darker curves correspond to smaller system sizes ($L=32$), while lighter curves represent larger systems ($L=256$). A clear narrowing effect is observed as $L$ increases.
To quantify this convergence, we performed a Finite-Size Scaling (FSS) analysis of the spectral width. Figure 4 demonstrates that the average singularity exponent $\langle\alpha\rangle$ converges asymptotically to the theoretical Ising value:
$\lim_{L\to\infty}\langle\alpha\rangle_{L}=0.875\pm 0.002.$ (17)
Crucially, the spectral width $\Delta\alpha$ scales as a power law of the system size, $\Delta\alpha\sim L^{-\gamma}$, vanishing in the thermodynamic limit. This collapse provides strong evidence that the multifractality observed in the pure Ising model is most likely a finite-size artifact largely influenced by the mixing of irrelevant operators and discreteness effects at small scales.
Figure 4: Finite-Size Scaling (FSS) of the singularity exponent $\langle\alpha\rangle$. The convergence towards the theoretical value $\alpha=0.875$ (green line) is evident as the system size increases ($1/L\to 0$).
IV.2 Probing Disorder: The RBIM as a Control
To demonstrate that the observed collapse to monofractality is not due to a lack of sensitivity in our method, we contrast these findings with the Random Bond Ising Model (RBIM). The RBIM represents a distinct universality class where quenched disorder is relevant (by the Harris criterion [11]), leading to the formation of rare Griffiths phases that support a true spectrum of local correlation lengths.
Figure 5 compares the singularity spectrum of the pure model against the disordered model at the same system size. The contrast is stark:
  • Pure Model: The spectrum collapses to a vertical line ($\Delta\alpha\to 0$), indicative of a single scaling dimension.
  • Disordered Model (RBIM): The spectrum remains broad and robust ($\Delta\alpha\approx 0.23$), persisting even after finite-size scaling.
This persistence confirms that our MFDFA protocol is sensitive enough to detect genuine multifractality when it exists. The ”dirty” criticality of the RBIM generates a legitimate hierarchy of exponents due to the spatial heterogeneity of the bond disorder, whereas the ”clean” criticality of the pure model is strictly monofractal. Consequently, the MFDFA spectral width $\Delta\alpha$ act as an effective diagnostic indicator for distinguishing disorder-dominated phases from pure critical fixed points.
Figure 5: Sensitivity of MFDFA to disorder. While the pure Ising model converges to a vertical line (green dashed line, $\Delta\alpha\to 0$), the Random Bond Ising Model (RBIM, red circles) [12] displays a broad multifractal spectrum ($\Delta\alpha\approx 0.23$), revealing the complex fluctuation structure induced by quenched impurities.
@@ V.1 Dependence on the Order of Detrending
V.2 Stability with Respect to the Scaling Range
Another critical aspect is the selection of the scaling window $s\in[s_{\min},s_{\max}]$ used to extract the exponent. To evaluate stability, we systematically varied the fitting bounds for the $L=256$ system, testing ranges from short-scale focused ($s\in[8,40]$) to broad-scale ($s\in[12,80]$).
We observe that the extracted Hurst exponent is highly stable, with a relative variation of approximately $1.9\%$ across all tested ranges. This stability confirms the existence of a robust, scale-invariant regime at criticality, distinct from microscopic lattice artifacts or finite-size boundary effects.
V.3 Shuffling Tests
To further validate that the measured correlations are intrinsic to the critical spin configurations and not a trivial consequence of the marginal spin distribution, we performed shuffling tests. Specifically, we randomly shuffled the spin values along each one-dimensional cut, thereby destroying spatial correlations while preserving the single-site distribution $\sigma\in\{-1,1\}$.
For sufficiently large system sizes ($L\geq 128$) where the scaling range allows for robust fitting, the shuffled profiles consistently yield a Hurst exponent consistent with the random walk limit ($H=0.5$). For our largest system ($L=256$), we obtained:
$H_{\mathrm{shuffled}}=0.498\pm 0.07.$ (19)
The stark contrast between this null result and the critical value observed in the original data ($H\approx 0.88$) confirms that the MFDFA signal originates from genuine long-range spatial correlations and not from finite-amplitude effects or probability density artifacts.
V.4 Comparison with Two-Point Correlation Functions
@@
V.5 Relation to Cluster-Based Observables

### 30 Mar 2026 v2

@@ Abstract
Multifractal Detrended Fluctuation Analysis (MFDFA) has emerged as a standard tool for characterizing scale invariance in complex systems, yet its application to discrete spin models is frequently marred by reports of “spurious multifractality” that contradict established theory. In this work, we resolve this controversy by establishing a rigorous protocol for the analysis of discrete lattice snapshots. Using the 2D Ising model as a benchmark, we demonstrate that the previously reported broad singularity spectra are finite-size artifacts dominated by lattice discreteness effects in the negative moment regime ($q<0$). By restricting the analysis to positive moments and performing a systematic Finite-Size Scaling (FSS) analysis, we show that the spectral width collapses to zero ($\Delta\alpha\to 0$) in the thermodynamic limit. The method accurately recovers the monofractal exponent of the Ising universality class ($\alpha\approx H\approx 0.875$), consistent with Conformal Field Theory. Applying this FSS protocol to the Random Bond Ising Model (RBIM), we reveal that the broad multifractal dispersion often attributed to quenched disorder also flattens systematically with increasing system size. These results demonstrate that apparent multifractality in finite discrete lattices is predominantly driven by pre-asymptotic discreteness artifacts, underscoring the universal necessity of FSS to avoid false positives in statistical mechanics.
@@ I Introduction
Several numerical studies have reported broad multifractal spectra in the two–dimensional Ising model and related variants [16, 9, 13, 17, 3]. These findings imply the existence of a hierarchy of scaling exponents, a result that contradicts the unique scaling dimension mandated by the RG fixed point. This discrepancy—often termed ”spurious multifractality”—raises a fundamental question: Does the reported multifractality reflect hidden complexity in the Ising phase transition, or does it signal a breakdown of the MFDFA methodology when applied to discrete variables in finite systems?
@@
To further probe the universality of these finite-size artifacts, we contrast our pure model findings with the Random Bond Ising Model (RBIM), characterized by quenched bond disorder and Griffiths phases. Strikingly, our FSS analysis reveals that the broad multifractal dispersion often reported for disordered lattices is also not robust. As the system size increases, the generalized Hurst exponent $H(q)$ systematically flattens. This establishes that even in systems with genuine structural heterogeneity, apparent multifractality at finite scales is heavily dominated by discreteness artifacts, making our FSS protocol universally mandatory.
@@ III.1 Model Hamiltonians and Monte Carlo Scheme
To ensure rigorous statistical convergence and eliminate the unsystematic noise typical of pre-asymptotic finite-size scaling, extensive computational resources were employed. For each system size, ranging from $L=32$ to $L=1024$, the lattice was thermalized using dynamic adjustments of Monte Carlo steps to overcome critical slowing down. For each temperature and system size, we collected an ensemble of $10^{4}$ independent equilibrium snapshots, separated by sufficient sweeps to guarantee statistical independence. From each 2D snapshot, a representative sample of up to 64 independent 1D spatial cuts was extracted and analyzed.
@@ 4. Scaling and Singularity Spectrum:
The MFDFA implementation employed here follows the standard algorithm established by Kantelhardt et al. [11], with optimizations for computational efficiency when analyzing large ensembles of lattice configurations. Our research group has successfully applied MFDFA and related multifractal methods to diverse systems including climate indices [2, 8], financial time series [1, 10], and biological related sequences [4, 12], establishing the methodological foundation for the present investigation of discrete lattice models.
@@ III.3 The Discrete Limit and Physical Cutoffs ($q>0$)
III.4 Mathematical Bounds: Divergence of Negative Moments and Scale Selection
The application of MFDFA to strictly discrete structural models, such as the Ising model ($\sigma_{i}\in\{-1,+1\}$), requires careful restrictions on the permissible moment range $q$ and fitting scales $s$ to avoid severe numerical artifacts.
In standard continuous time series, negative moments ($q<0$) are utilized to amplify the scaling behavior of small fluctuations. However, in binary discrete systems, the local variance $F^{2}(s,v)$ within a segment of length $s$ can be exactly zero if the segment falls entirely within a “frozen” magnetic domain (where all spins are identical). According to Eq. 12, the fluctuation function for negative moments depends on $F^{2}(s,v)]^{-|q|/2}$. Consequently, any segment with zero local variance will cause the summation to mathematically diverge to infinity. This singularity completely dominates the fluctuation function, creating an artificially broad, “spurious” singularity spectrum. Therefore, to evaluate the genuine critical scaling of the domains rather than zero-variance division artifacts, the analysis must be strictly bounded to positive moments ($q>0$). In our revised analysis, we robustly evaluate the positive regime to capture the true thermodynamic correlations.
Another critical aspect is the selection of the scaling window used to extract the exponent. To ensure uniformity and eliminate heuristic biases, all scaling exponents and singularity spectra presented in this work were extracted using the standard, dynamically adjusted fitting bounds: $s\in[10,L/4]$. The lower bound $s_{min}=10$ prevents the polynomial detrending from overfitting the microscopic lattice spacing. The upepr cutoff $s_{max}=\lfloor L/4\rfloor$ ensures that a minimum of four independent, non-overlapping macroscopic segments are available per profile, which is the statistical minimum required to perform a reliable spatial average for the fluctuation functions.
@@ IV.1 The Critical Fixed Point: Restoration of Monofractality
We first examine the scaling behavior of the pure 2D Ising model. A prerequisite for any valid multifractal analysis is the accurate recovery of the known Hurst exponent at the critical point.
@@
To rigorously test the accurate recovery of the known Hurst exponent at the critical point, we expanded our simulations to an extensively large lattice size of $L=1024$, significantly increasing the Monte Carlo thermalization steps to ensure absolute equilibrium. Figure 2 displays the temperature dependence of the second-order Hurst exponent $H(q=2)$ for this system. As the temperature approaches the critical point $T_{c}\approx 2.269$, the exponent exhibits a sharply defined crossover, intersecting the theoretical prediction $H=0.875$ precisely at the phase transition. Furthermore, the high-temperature regime rapidly converges to the uncorrelated random-walk limit ($H=0.5$). This high-precision agreement unequivocally confirms that our spatial MFDFA protocol correctly captures the anomalous dimension of the magnetization field, satisfying the theoretical scaling relation $H=1-\eta/2$ without finite-size distortion at this scale.
@@
Figure 2: Evolution of the Hurst exponent $H$ ($q=2$) vs Temperature for $L=1024$. The intersection with the theoretical value $H=0.875$ (blue dotted line) occurs precisely at the critical temperature $T_{c}$ (red dashed line).
@@
The central result of this study concerns the width of the singularity spectrum $f(\alpha)$. Previous studies reporting “spurious multifractality” typically observed a broad spectral width. In Figure 3, we present the robust evolution of the singularity spectrum for system sizes ranging up to $L=1024$, strictly in the positive moment regime ($q>0$). As $L$ increases, we observe a dramatic narrowing of the spectrum. The distinct branches progressively collapse onto a single vertical line centered identically at the theoretical limit $\alpha\approx 0.875$. This visual evidence strongly suggests that the broad spectra observed at smaller sizes ($L\leq 128$) are entirely finite-size artifacts driven by lattice discreteness, rather than an intrinsic multifractal property.
Figure 3: Evolution of the singularity spectra $f(\alpha)$ for $q>0$ and system sizes from $L=32$ to $L=1024$. As the system size increases the spectra exhibit a highly systematic narrowing and a monotonic shift toward the theoretical prediction. The macroscopic lattice ($L=1024$) collapses into a near-vertical line precisely at the theoretical monofractal dimension $\alpha=0.875$ (dashed green line). Error bars (standard error of the mean) are smaller than the symbol size due to ensemble averaging.
To rigorously quantify this convergence we performed a Finite-Size Scaling (FSS) analysis of the spectral width. Figure 4a shows the Finite-Size Scaling (FSS) of the central singularity exponent $\alpha$ against the inverse system size $1/L$. While finite-size effects and lattice cutoffs induce a non-monotonic behavior at intermediate scales, the asymptotic behavior in the thermodynamic limit ($1/L\rightarrow 0$) shows a convergence towards the exact theoretical Ising value of $\alpha=0.875$.
To explicitly demonstrate that the broad spectral width is a finite-size artifact, Figure 4b plots $\Delta\alpha$ as a function of the system size $L$ on a logarithmic scale. We exclude the smallest lattice ($L=32$) from the power-law fit, as its microscopic scale falls within the pre-asymptotic regime where MFDFA detrending windows are too restricted to capture true scaling. Within the asymptotic regime ($L\geq 64$), the spectral width exhibits a power-law decay, $\Delta\alpha\sim L^{-0.57}$. This finite-size scaling confirms that the apparent multifractality systematically vanishes in the thermodynamic limit ($L\rightarrow\infty$), leaving a monofractal singularity.
Figure 4: Finite-Size Scaling (FSS) analysis of the spurious multifractality. (a) Convergence of the central singularity exponent $\alpha$ as a function of the inverse system size $1/L$. The asymptotic limit strictly recovers the theoretical 2D Ising prediction $\alpha=0.875$ (dashed green line). (b) Log-log plot of the spectral width $\Delta\alpha$ versus system size $L$ for the positive moment regime ($q>0$). Excluding the pre-asymptotic microscopic lattice ($L=32$), the data exhibits a robust power-law decay $\Delta\alpha\sim L^{-0.57}$ (dashed red line), confirming that the multifractal spectrum collapses into a single dimension in the thermodynamic limit.
IV.2 Finite-Size Scaling of the Random Bond Ising Model
To investigate whether the severe finite-size distortion observed in the pure model also affects systems with heterogeneous spatial correlations, we applied our MFDFA Finite-Size Scaling protocol to the Random Bond Ising Model (RBIM). The RBIM introduces quenched bond disorder, leading to the formation of rare Griffiths phases that order locally at temperatures deviating from the bulk $T_{c}$.
Figure 5 presents the generalized Hurst exponent $H(q)$ for the RBIM across multiple system sizes ($L\in[64,1024]$), evaluated in the robust positive moment regime. While small disordered lattices ($L=64$) exhibit a broad dispersion curve ($\Delta H\approx 0.825$), our FSS analysis reveals that this width is not static, but decreases with system size. However, the critical insight lies not merely in the presence of a collapse, but in its rate.
Figure 5: Finite-size scaling of the generalized Hurst exponent $H(q)$ for the Random Bond Ising Model (RBIM) evaluated in the physically robust positive moment regime ($q>0$). In a strictly monofractal system, $H(q)$ is a constant horizontal line. While small disordered lattices ($L=64$) exhibit significant dispersion ($\Delta H\approx 0.825$), the $H(q)$ curves systematically flatten as the system size increases up to $L=1024$ ($\Delta H\approx 0.196$), progressively approaching the asymptotic monofractal limit ($H=0.875$, dashed green line). The rate of this flattening is fundamentally distinct from the pure Ising model, serving as a signature of quenched disorder.
In the pure Ising model, the pre-asymptotic discreteness artifacts decay rapidly, following a steep power law ($\Delta\alpha\sim L^{-0.57}$) that unequivocally drives the spectral width to zero in the thermodynamic limit. In contrast, the RBIM exhibits a systematically much slower collapse. As illustrated in Figure 6, the progression up to the macroscopic scale of $L=1024$ shows that the decay rate of $\Delta H$ decelerates significantly. An asymptotic fit of the form $\Delta H\sim AL^{-\gamma}+\Delta H_{\infty}$ captures this saturation, revealing a strictly non-zero residual dispersion of $\Delta H_{\infty}\approx 0.123$ in the thermodynamic limit ($L\to\infty$).
This quantitative difference in the FSS collapse rate constitutes a fundamental physical distinction between the two universality classes. In the pure model, the spectral width originates entirely from pre-asymptotic discreteness effects and vanishes in the thermodynamic limit. In the RBIM, by contrast, the quenched impurities generate rare Griffiths regions with locally diverging correlation lengths, producing a genuine residual multifractality ($\Delta H_{\infty}\approx 0.123$) that survives the scaling limit. Consequently, our FSS protocol serves a dual and complementary purpose: it eliminates rapid pre-asymptotic discrete artifacts common to both models, while simultaneously revealing genuine disorder-dominated complexity through the persistence of a finite asymptotic dispersion — a feature absent in clean criticality.
Figure 6: Comparison of the Finite-Size Scaling (FSS) collapse rates between pure and disorder-dominated criticality. In the pure 2D Ising model (black squares), the multifractal dispersion $\Delta\alpha$ decays rapidly as a steep power law ($\gamma\approx 0.57$), strictly vanishing in the thermodynamic limit ($L\to\infty$). Conversely, the Random Bond Ising Model (RBIM, purple circles) exhibits a systematically slower decay of its generalized Hurst dispersion $\Delta H$. An asymptotic fit (dashed purple line) extending up to $L=1024$ reveals a significant saturation trend, firmly establishing a non-zero residual multifractality of $\Delta H_{\infty}\approx 0.123$ in the macroscopic limit. This quantitative divergence in collapse rates allows MFDFA to effectively discriminate between transient discreteness artifacts and genuine Griffiths phase complexity.
@@ V.1 Dependence on the Order of Detrending
V.2 Shuffling Tests
TTo definitively rule out the possibility that the observed scaling behaviors are mere artifacts of the binary probability density function of the discrete spins , we performed a surrogate data analysis using a rigorous shuffling test. If the scaling extracted by MFDFA is genuinely driven by the topological spatial correlations of the critical clusters, randomly shuffling the spatial cuts should destroy this memory, reducing the system to uncorrelated white noise.
To capture the full extent of this effect, we evaluated the full generalized Hurst exponent H(q) over the structurally robust positive moment regime for the macroscopic lattice ($L=1024$). Figure 7 illustrates the profound impact of the shuffling procedure. While the original spatial cuts exhibit a non-trivial dispersion curve approaching the theoretical Ising limit (H=0.875), the randomly shuffled surrogate data undergoes a complete systemic collapse. The $H(q)$ curve drops abruptly to a flat, horizontal line at $H(q)\approx 0.5$ across all evaluated moments.
This perfect flattening indicates a strict generalized width of $\Delta H\approx 0$, which mathematically translates to a singularity spectrum width of $\Delta\alpha\rightarrow 0$. The slight deviation from the exact theoretical $0.5$ value is a standard, well-documented finite-size artifact associated with the polynomial detrending in MFDFA. This full-spectrum null result confirms that the critical scaling captured by our protocol is purely structural and topologic, verifying that MFDFA correctly identifies the absence of correlation when the physical domain structure is artificially destroyed.
Figure 7: Surrogate data analysis (shuffling test) for the macroscopic lattice ($L=1024$) at the critical temperature, visualized via the generalized Hurst exponent $H(q)$. Randomly shuffling the spatial cuts destroys all topological correlations while preserving the single-site probability density. Consequently, the critical scaling behavior (red squares, near the exact Ising limit $H=0.875$) abruptly collapses to a flat, uncorrelated line representing standard white noise (blue circles, $H\approx 0.49\pm 0.01$). The slight deviation from the theoretical $H=0.5$ (dashed gray line) is a standard finite-size artifact of the MFDFA polynomial detrending. This null result unambiguously confirms that the scaling observed in the original cuts is structural and genuine.
V.3 Comparison with Two-Point Correlation Functions
@@
V.4 Relation to Cluster-Based Observables
@@
V.5 Explicit Distribution of Local Hölder Exponents
As suggested by recent methodological discussions comparing MFDFA with continuous wavelet-based formalisms (e.g., WTMM), an explicit point-wise evaluation of the local scaling properties provides a powerful complementary validation. While WTMM extracts local singularities via the continuous wavelet transform, a mathematically equivalent pointwise estimation of the local Hölder exponents, $h(x)$, can be achieved through a localized sliding-window scaling analysis. This approach maintains methodological consistency with our fluctuation-based framework while avoiding the severe boundary artifacts that wavelets often encounter when applied to strictly discrete, step-like binary data.
To corroborate our MFDFA findings, we conducted an explicit study of the local Hölder exponents along the spatial cuts of our $L=1024$ system at the critical temperature. We employed a localized sliding-window scaling analysis with a window size of $W=256$. As illustrated in Figure 8, the resulting probability density function of the local Hölder exponents is highly localized. It forms a narrow, Gaussian-like distribution ($\sigma\approx 0.15$) centered at $h>\approx 0.91$ , in excellent agreement with the exact theoretical prediction of 0.875 (with the slight positive deviation being a standard finite-window effect). Crucially, the distribution lacks the broadness, heavy tails, or strong asymmetry characteristic of true multifractal cascades. This explicit pointwise measurement definitively confirms that the critical 2D Ising model is governed by a single dominant scaling exponent, mathematically corroborating the collapse of the spectral width observed in our MFDFA finite-size scaling protocol.
Figure 8: Explicit probability density function of the local Hölder exponents, $h(x)$, for the macroscopic 2D Ising model ($L=1024$) at the critical temperature. The exponents were extracted pointwise using a localized sliding-window scaling analysis. The resulting histogram forms a narrow, highly localized Gaussian distribution ($\sigma\approx 0.154$), lacking the broadness or heavy tails characteristic of genuine multifractal cascades. The empirical mean ($\langle h\rangle\approx 0.908$, dotted red line) is in excellent agreement with the exact theoretical prediction of $\alpha=0.875$ (dashed green line), with the slight positive deviation being a standard finite-window boundary effect. This explicit local measurement definitively confirms that the system is governed by a single dominant scaling exponent, mathematically corroborating the monofractal collapse observed in our MFDFA finite-size scaling protocol.